Can you solve this for me? Fnd the slope, delta y, delta x, y intercept, x intercept, midsection, and distance given only this information:
(-6,14) (18,-4)

Answer by stanbon(75887)

Fnd the slope, delta y, delta x, y intercept, x intercept, midsection, and distance given only this information:
(-6,14) (18,-4)
---------------------
delta y = [14--4]=18
delta x = [-6-18]=-24
slope = 18/-24 = -3/4
--------------------------
midpoint = [(-6+18)/2,(14+-4)/2] = (6,5)
----------------------------
distance = sqrt[18^2+(-24)^2] = sqrt(900) = 30
-------------------------------
Equation has form y=mx+b; need to find "b".
14=(-3/4)(-6)+b/
14=(9/2)+b
b = (28-9)/2 = 19/2
EQUATION: y = (-3/4)x + 19/2
----------------
y-intercept:
Let x=0 then y = 19/2
-------------------
x-intercept:
Let y=0, then (-3/4)x+19/2 = 0 ; x = (19/2)(4/3)= 38/3
============================================================
Cheers,
Stan H.